Operations Report March 29th

Crew 191 Operations Report 29/03/2018

SOL:5

Name of person filing report: Wataru Okamoto

Non-nominal systems: Solar System

Notes on non-nominal systems: Solar system seems unusual in this morning. Now it looks stable.

Generator (hours run):13hours

Solar SOC 93% (Before generator is run at night)

Diesel under 25%

Propane 72psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) ? 7gallons.

Water (trailer) 60%

Water (static) 70%

Trailer to Static Pump used ? no

Water (loft) Static to Loft Pump used yes

Water Meter: 16:00 132835.5 

Toilet tank emptied: no

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) –

Oil Added –

ATV Fuel Used: N/A

Hours the ATVs were Used today: N/A

Notes on ATVs: N/A

Deimos rover used: –

Hours:116.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: 100%

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: –

Hours:27.9h

Beginning charge:100%

Ending charge: 100%

Currently charging: 100%

Opportunity rover used: Morisawa, Takeda

Hours: 27.4h

Beginning charge: 100%

Ending charge: 99%

Currently charging: 99%

Curiosity rover used: Tsukishiro

Hours:22.7h

Beginning charge: 100%

Ending charge: 97%

Currently charging: 97%

HabCar used and why, where? –

General notes and comments: Solar system seems unusual in this morning. We have discussion about this problem with Dr. Shannon and now it seems stable. We will keep monitoring.

Summary of internet: Under control

Summary of suits and radios: The surface of helmet polished by Plexus & microfiber clothes.

Summary of Hab operations: Shower heater doesn’t work.

Summary of GreenHab operations: No problem

Summary of ScienceDome operations: Only use checking the solar system.

Summary of RAMM operations: N/A

Summary of health and safety issues: No problem

Questions, concerns and requests to Mission Support:

The PLEXUS splay is almost empty. We need another supply.

We will put some garbage bags into the engineering airlock tonight. Please take them out.